# Chained computerized adaptive testing for the Force Concept Inventory

### v1

@@ III.1 Typical numerical simulation result
Figure 2 shows the simulation results assuming a linear model for the true proficiency progression . The test length of each administration was set to . The standard deviation of the prior distribution was fixed at . The left figure of Fig. 2 is a result without collateral information (basic CAT) and the right figure is a result with collateral information. The graphs include the simulation results for nine-time points from to , where the true proficiency level increases from to (see the purple lines in the graphs). For each time point, we generated 10 000 responses based on the same true theta at a given moment in time and estimated the proficiency level for each response, then we calculated the average and the standard deviation of the estimated theta (the yellow and blue lines). The standard deviation for the estimates of the mean is about 0.0025. The bias was calculated from the difference between the true theta and the average of the estimated theta (the red lines), then the RMSE was calculated using Eq. (2) (the green lines). In Fig. 2 (b), the standard deviation is much larger than the bias, so the SD values and the RMSE values are very close. In the graphs, the reference values are shown at and on the horizontal axis, where the statistics were calculated using the true proficiency levels at and and administering the whole test items (30 FCI items) in both cases without collateral information.
@@ III.2 Comparing overall trends
As we mentioned above, the standard deviation of the prior distribution is a free parameter. We found that if is small, the variance in Eq. (4) tends to be small, but the bias in Eq. (3) tends to be large when the proficiency progression is large. If is large, the variance tends to be large, but the bias tends to be relatively small when the proficiency progression is large. Because we do not know the amount of proficiency progression before the observation, we set the value of such that the RMSE is minimized on average. Specifically, for a given value of (=0.1, 0.2, …, 1.0), we conducted simulations and calculated the time-averaged RMSE in Eq. (5) from to . For the pre-post method, the time-averaged RMSE was calculated for and . Then we calculated the overall average in Eq. (6) for all progression models. After that, we chose the value of which minimizes the overall average of the RMSE. We obtained the following values for each case, namely, for the case without collateral information, for the case with collateral information, and for the pre-post method.
@@
After fixing the value of which is most likely to give the highest efficiency for each case, we calculated the time-averaged RMSE. Table 1 shows the result for . In Table 1, we can observe that the time-averaged RMSE in the case with collateral information is smaller than or comparable to that in the case without collateral information for all progression models. Moreover, the time-averaged RMSE in the case with collateral information is smaller than that of the pre-post case for most of the progression models. Especially, the case with collateral information is much more efficient when the absolute value of at final is large as in the Constant model (1) and the Linear model (). On the other hand, the case without collateral information has a comparable efficiency in the Linear model (.5) and the Step model (), where the absolute value of varies near zero and the absolute value of at final is small. One of the causes of this could be that the case without collateral information always assumes the mean value of prior distribution . Finally, the pre-post case is more efficient when the absolute value of at initial is large and at final is close to 0 as in the Linear model (). This is because for CAT with short test length, such as , if the absolute value of at is large, the RMSE at is much larger than that of the pretest where .
@@ III.3 Length dependence, content balancing, and item exposure control
Table 2 shows the dependence of the overall average RMSE in Eq. (6) of Chain-CAT (with collateral information) on the test length of each administration. The standard deviation of the prior distribution was fixed at the optimal value, . Table 2 also includes a comparison of the cases with/without content balancing and item exposure control. In Table 2, we can see that the overall average RMSE increases as the test length decreases. However, the overall averaged RMSE of Chain-CAT is comparable to that of the pre-post method even when the test length is (compare the value of 0.275 in Table 1 to that of 0.277 in Table 2). This means that if we administer the Chain-CAT nine times in a course, by using the collateral information, we could reduce the total test length in a course to 45 items from the pre-post method (60 items) without compromising the test accuracy and precision if we are not concerned about improving the validity of the test. On the other hand, by implementing content balancing and item exposure control, the overall average RMSE is increased from 0.277 to 0.317 when , an increase of about 10%. This is because when the choice of administration is limited, the items that result in lower accuracy and precision are administered. When , the overall average RMSE with content balancing and item exposure control is smaller than that of the pre-post method (compare the value of 0.275 in Table 1 to that of 0.269 in Table 2). In this case, the total test length is 90 items which is longer than that of the pre-post method, but the accuracy and precision are increased.
@@
Figure 3 shows histograms of how many times an item was administered in the nine tests when in the Linear model (). The horizontal axis shows the number of times an item was administered, and the vertical axis shows the frequency. The left (right) graph shows the result without (with) content balancing and item exposure control. Recall that our simulation had 10 000 simulees take nine tests. In the case of , this totals 450 000 items administered in the simulation. If we count any time that any of the 30 FCI items showed up in all nine of a simulee’s tests, we find that occurs many times when we do not employ content balancing and item exposure control (right-most bar of the left graph in Fig. 3). We can also see that there are even more cases where an item was never administered (left-most bar of the left graph in Fig. 3) in that semester.
@@
The implementation of content balancing and item exposure control changes this situation, as shown in the right graph of Fig. 3. First, for content balancing, we ensured that the items in each test are not biased toward specific concepts. For example, since there are five subgroups, there was one item from each subgroup on each test when . In addition to this, following the calculation in Sec. II.5.3, we set as the maximum number of times each item could be administered in the nine tests for . We can see that these are realized in the right graph in Fig. 3. However, even with content balancing and item exposure control, there are many cases where an item is never administered in that semester. This problem is evident in the bottom graph in Fig. 4. Compared to the top graph, the number of items with non-zero percentages has increased. However, nearly half of the items with low discrimination were never administered. (Although Q.10 and Q.12 have rather high discrimination, they were almost never administered due to their very small difficulty parameters: less than .)
These results show that simply expanding the item bank may be insufficient to improve the situation further. To improve the test efficiency, we suggest creating or incorporating items from other assessment tests [48, 49, 50, 51] that are highly discriminative, i.e. that can better distinguish respondents with different levels of proficiency.
@@ IV.1 Summary
To address the issues of the pre- and post-paradigm, we designed a Chain-CAT algorithm that sequentially links the results of each CAT administration in a course using collateral information. We analyzed the efficiency of the Chain-CAT algorithm through numerical simulations of various progression models for the values of true proficiency. The simulation results showed that the Chain-CAT algorithm significantly improved the test efficiency in most of the progression models compared to the simple CAT without collateral information. The simulation results indicated that the accuracy and precision of the Chain-CAT are comparable to that of the pre-post method even when each test length is as short as 5 items when the test is administered 9 times; therefore, the total test length of Chain-CAT can be less than that of the pre-post method. This finding, however, is without imposing constraints of item balancing and item exposure control. When these constraints are imposed, we found that the efficiency of the Chain-CAT is lower than that of the pre-post method when the length of each of the nine tests is five items. These results suggest that the item bank should be expanded with new items or by incorporating other inventory items with high discrimination.
@@ IV.2 Limitations and future work
The Chain-CAT can be validated and improved by means other than content balancing and item exposure control. First, although we found that our FCI dataset has sufficient unidimensionality and local independence at the overall test level [40, 41], higher unidimensionality and local independence are desirable. This can be achieved by using multidimensional models [52, 53, 54, 55] and dropping one item from each locally dependent pair. Adding new items would be necessary to compensate for the reduction in efficiency.
Second, it is meaningful to conduct Chain-CAT on real classes and assess the validity of the instrument. We can analyze the validity of the real class data by analyzing the consistency between the Chain-CAT results and other assessment data, such as that obtained in clinical interviews. We have conducted a pilot study for validation where we deployed the Chain-CAT system using the mirtCAT package [56] with shinyapps. Students used their smartphones to complete the Chain-CAT. The preliminary results showed that Chain-CAT results are roughly consistent with interview findings. Further work should investigate this more extensively.
@@
In addition to the methods we used in this study, there are many options to explore that may further improve the efficiency of the Chain-CAT. Although we used the length criterion as the stopping step, an alternative option is to use the precision criterion. This criterion administers as many items as necessary until pre-specified precision is obtained [20]. However, because the test length varies from respondent to respondent when the precision criterion is used, it would be inconvenient to administer the test in a classroom setting, so at-home administration may be preferable.
The algorithm for the use of collateral information could be improved. In the test step of the Chain-CAT algorithm, we directly used a given student’s proficiency estimate; namely, we chose the prior distribution as a normal distribution with for the second test. The other possibility is predicting the initial proficiency estimate of the th test by a regression model in a similar way to [31] or another time-series model using the given student’s proficiency estimates before the th test. Another option is to implement the algorithm utilizing the response time in addition to the response choice of the students [57]. This information might further improve the test efficiency.
@@
The following issues are needed to develop the feedback system based on Chain-CAT FCI. First, it is necessary to develop a method to analyze the progression of individual proficiency. Our simulation results showed a clear progression trend as shown in Fig. 2, because the statistics were calculated using 10 000 simulated responses at each time point. However, for an individual respondent, the progression of proficiency fluctuates much more, which makes it difficult to understand the state of the student precisely. Therefore, we may need to estimate a student’s state based on the analysis of multiple test results, for example, by calculating a time average. It might be appropriate to group the proficiency estimates into several levels as in [33].
There are many ways to provide feedback from Chain-CAT to respondents [58]. A simple score, the proficiency estimate would not be useful for students. It would be more useful, however, to group the proficiency estimates into levels, for example, simply dividing into four levels by setting thresholds. In Chain-CAT, students can receive feedback in the form of graphs showing their progress, as well as in the form of sentences with encouragement and/or guidance provided by generative artificial intelligence [59]. Ideally, this feedback will promote reflection, metacognition, and self-regulation for the students [60]. These tasks would be our future work.

### v2

@@ III.1 Typical numerical simulation result
Figure 2 shows the simulation results assuming a linear model for the true proficiency progression . The test length of each administration was set to . The standard deviation of the prior distribution was fixed at . The left figure of Fig. 2 is a result without collateral information (basic CAT) and the right figure is a result with collateral information. The graphs include the simulation results for nine-sessions from to , where the true proficiency level increases from to (see the purple lines in the graphs). For each session, we generated 10 000 responses based on the same true theta at a given session and estimated the proficiency level for each response, then we calculated the average and the standard deviation of the estimated theta (the yellow and blue lines). The standard deviation for the estimates of the mean is about 0.0025. The bias was calculated from the difference between the true theta and the average of the estimated theta (the red lines), then the RMSE was calculated using Eq. (2) (the green lines). In Fig. 2 (b), the standard deviation is much larger than the bias, so the SD values and the RMSE values are very close. In the graphs, the reference values are shown at and on the horizontal axis, where the statistics were calculated using the true proficiency levels at and and administering the whole test items (30 FCI items) in both cases without collateral information.
@@ III.2 Comparing overall trends
As we mentioned above, the standard deviation of the prior distribution is a free parameter. We found that if is small, the variance in Eq. (4) tends to be small, but the bias in Eq. (3) tends to be large when the proficiency progression is large. If is large, the variance tends to be large, but the bias tends to be relatively small when the proficiency progression is large. Moreover, we found that the rank order of the overall average RMSE magnitude depends on the value of . For example, when is set to 1 for all methods, the pre-post method shows the highest accuracy (lowest RMSE), whereas when is set to 0.3, the Chain-CAT with collateral information shows the highest accuracy, and so on. We were concerned that comparing results at a fixed could introduce arbitrariness, depending on the analyst’s choice of the common value. Therefore, we calculated the overall average across various values and based our comparison on the that yielded the lowest overall average RMSE. This approach reflects practical considerations for implementing the test in classroom settings, where should be determined before administration without knowing the amount of proficiency progression in advance. In this case, it is reasonable to choose that yielded the lowest overall average in each method.
Specifically, for a given value of (=0.1, 0.2, …, 1.0), we conducted simulations and calculated the session-averaged RMSE in Eq. (5) from to . For the pre-post method, the session-averaged RMSE was calculated for and . Then we calculated the overall average in Eq. (6) for all progression models that we assumed in the simulations. After that, we chose the value of which minimizes the overall average of the RMSE. We obtained the following values for each case, namely, for the case without collateral information, for the case with collateral information, and for the pre-post method.
@@
After fixing the value of which is most likely to give the highest efficiency for each case, we calculated the session-averaged RMSE. Table 1 illustrates the effect of collateral information within each progression model for the case of . For each model, we compare the effectiveness of different methods: without CI, with CI, and the pre-post approach. To isolate and better understand the impact of collateral information, we did not incorporate content balancing or item exposure control in this table. In Table 1, we can observe that the session-averaged RMSE in the case with collateral information is smaller than or comparable to that in the case without collateral information for all progression models. Moreover, the session-averaged RMSE in the case with collateral information is smaller than that of the pre-post case for most of the progression models. Especially, the case with collateral information is much more efficient when the absolute value of at final is large as in the Constant model (1) and the Linear model (). On the other hand, the case without collateral information has a comparable efficiency in the Linear model (.5) and the Step model (), where the absolute value of varies near zero and the absolute value of at final is small. One of the causes of this could be that the case without collateral information always assumes the mean value of prior distribution . Finally, the pre-post case is more efficient when the absolute value of at initial is large and at final is close to 0 as in the Linear model (). This is because for CAT with short test length, such as , if the absolute value of at is large, the RMSE at is much larger than that of the pretest where .
@@ III.3 Length dependence, content balancing, and item exposure control
Table 2 shows the dependence of the overall average RMSE of Chain-CAT (with collateral information) on the test length of each administration. The overall average RMSE is calculated by taking the average of the session-averaged RMSEs across all progression models assumed in the simulations, as defined in Eq. (6). The standard deviation of the prior distribution was fixed at the optimal value, . Table 2 also includes a comparison of the cases with/without content balancing and item exposure control. In Table 2, we can see that the overall average RMSE increases as the test length decreases. However, the overall averaged RMSE of Chain-CAT is comparable to that of the pre-post method even when the test length is (compare the value of 0.275 in Table 1 to that of 0.277 in Table 2). This means that if we administer the Chain-CAT nine times in a course, by using the collateral information, we could reduce the total test length in a course to 45 items from the pre-post method (60 items) without compromising the test accuracy and precision if we are not concerned about improving the validity of the test. On the other hand, by implementing content balancing and item exposure control, the overall average RMSE is increased from 0.277 to 0.317 when , an increase of about 10%. This is because when the choice of administration is limited, the items that result in lower accuracy and precision are administered. When , the overall average RMSE with content balancing and item exposure control is smaller than that of the pre-post method (compare the value of 0.275 in Table 1 to the value of 0.269 in Table 2). In this case, the total test length is 90 items which is longer than that of the pre-post method, but the accuracy and precision are increased. This improvement is partly attributable to the assessment using 50% more items.
As noted above, applying content balancing or item exposure control limits the flexibility of item selection, resulting in the administration of items that reduce accuracy and precision. The findings above indicate that the current 30-item FCI is insufficient for implementing Chain-CAT effectively, and that expanding the item bank—either by developing new items or incorporating items from other concept inventories—is necessary. In this context, the results shown in Table 1 and Table 2, obtained without content balancing or item exposure controls, offer valuable insights. They can be interpreted as representing a future scenario in which the FCI item bank has been expanded to include a sufficient number of replacement items with similar difficulty and discrimination characteristics. In other words, from these findings, we expect that such a level of measurement accuracy could be attainable if the FCI item bank is expanded, even with content balancing and item exposure controlled.
@@
Figure 3 shows histograms of how many times an item was administered in the nine tests when in the Linear model (). The horizontal axis shows the number of times an item was administered, and the vertical axis shows the frequency. If an item appears in all nine tests, its frequency contributes one count to the bin labeled “9” in the histogram representing the number of times administrated. The left (right) graph shows the result without (with) content balancing and item exposure control. Recall that our simulation had 10 000 simulees take nine tests. In the case of , this totals 450 000 items administered in the simulation. In the histogram, the sum of the products of each value (number of times administrated) and its corresponding value (frequency) equals the total number of item administrations in the simulation, which is 450 000. If we count any time that any of the 30 FCI items showed up in all nine of a simulee’s tests, we find that occurs many times when we do not employ content balancing and item exposure control (right-most bar of the left graph in Fig. 3). We can also see that there are even more cases where an item is never administered (left-most bar of the left graph in Fig. 3) in that semester.
@@
The implementation of content balancing and item exposure control changes this situation, as shown in the right graph of Fig. 3. First, for content balancing, we ensured that the items in each test are not biased toward specific concepts. For example, since there are five subgroups, there was one item from each subgroup on each test when . In addition to this, following the calculation in Sec. II.5.3, we set as the maximum number of times each item could be administered in the nine tests for . We can see that these are realized in the right graph in Fig. 3. However, even with content balancing and item exposure control, there are many cases where an item is never administered in that semester. This problem is evident in the bottom graph in Fig. 4. Compared to the top graph, the number of items with non-zero percentages has increased. However, nearly half of the items with low discrimination were never administered. (Although Q.10 and Q.12 have rather high discrimination, they were almost never administered due to their very small difficulty parameters: less than .) One additional issue concerns the variability in difficulty parameters. Since three of the content areas (KN, SL, and TL) contain five or fewer items, the range of difficulty parameters is limited. For example, the difficulty parameters of the SL items are 0.471 for Q.21, for Q.22, and 0.567 for Q.26. This limited variability may be insufficient to accurately and precisely estimate students’ abilities.
These results show that simply expanding the item bank may be insufficient to improve the situation further. To improve the test efficiency, we suggest creating or incorporating items from other assessment tests [52, 53, 54, 55] that are highly discriminative, i.e. that can better distinguish respondents with different levels of proficiency. In addition, the items should display sufficient variability in their difficulty parameters.
@@ IV.1 Summary
To address the limitations of the pre-post method, we proposed an alternative approach using a Chain-CAT algorithm that sequentially links the results of each CAT administration in a course using collateral information. We analyzed the efficiency of the Chain-CAT algorithm through numerical simulations of various progression models for the values of true proficiency. The simulation results showed that the Chain-CAT algorithm significantly improved the test efficiency in most of the progression models compared to the simple CAT without collateral information. The simulation results indicated that the accuracy and precision of the Chain-CAT are comparable to that of the pre-post method even when each test length is as short as 5 items when the test is administered 9 times; therefore, the total test length of Chain-CAT can be less than that of the pre-post method. This finding, however, is without imposing constraints of item balancing and item exposure control. When these constraints are imposed, we found that the efficiency of the Chain-CAT is lower than that of the pre-post method when the length of each of the nine tests is five items. These results suggest that the item bank should be expanded by adding new items or incorporating items from other inventories that have high discrimination and exhibit sufficient variability in difficulty parameters.
@@ IV.2 Limitations and future work
There are many possibilities for further improving the validity of the Chain-CAT. First, although we found that our FCI dataset has sufficient unidimensionality and local independence at the overall test level [43, 44], higher unidimensionality and local independence are desirable. This can be realized by dropping one item from each locally dependent pair. Adding new items would be necessary to compensate for the reduction in efficiency. Alternatively, using multidimensional models may produce a better fit to the data, resulting in more valid proficiency estimates [56, 57, 58, 59].
Second, our content balancing algorithm utilized the FCI taxonomy table [48], but researchers have identified issues with this grouping system [54, 60, 61, 62, 63, 64, 65]. When new highly discriminative items (as mentioned above) are added to the item bank, the reliability of the statistical grouping and the content balancing algorithm should be examined.
Third, it is meaningful to conduct Chain-CAT on real classes and assess the validity of the instrument. We can analyze the validity of the real class data by analyzing the consistency between the Chain-CAT results and other assessment data, such as that obtained in clinical interviews. We have conducted a pilot study for validation where we deployed the Chain-CAT system using the mirtCAT package [66] with shinyapps. Students used their smartphones to complete the Chain-CAT. The preliminary results showed that Chain-CAT results are roughly consistent with interview findings. Moreover, the results showed that the progression of proficiency estimates was relatively stable and interpretable—for example, displaying either a steady linear increase or consistently low values—rather than highly fluctuating. Further work should investigate this more extensively.
Fourth, in Sec. III.2, we calculated the overall average RMSE across various values and based our comparison on the that yielded the lowest overall average RMSE. This approach reflects practical considerations for implementing the test in classroom settings, where instructors determine before administration without knowing the amount of proficiency progression in advance. A possible real-world usage scenario would be one in which the teacher chooses either Chain-CAT (with CI) or the pre-post method. In the former (latter) case, the program would automatically select =0.3 (=0.6) to minimize RMSE. We found that, even when sigma is optimized, the minimal RMSE when collateral information is not used is larger than the RMSE when collateral information is used. As such, we do not recommend using Chain-CAT without collateral information. This analysis using session and overall averages may not provide a rigorous comparison, but it can offer a useful reference from a practical perspective that considers the actual conditions of test administration in the classroom. Since our method of analysis is exploratory, there remains room for improvement in such comparative approaches.
Fifth, when expanding the item bank by developing new items, it is important to recognize that evaluating what constitutes a “good” question requires multiple perspectives. The results presented in Fig. 4 illustrate one method of item evaluation, grounded in item response theory. While some items appear to be more discriminating and therefore more informative from the IRT perspective, this approach captures only a portion of what defines item quality. Other important factors—beyond those identified by IRT—may be valued by experts. Thus, to meaningfully assess the quality of new items, additional investigation is needed to determine whether items that perform well under IRT are also regarded as high quality by experts.
@@
In addition to the methods we used in this study, there are many options to explore that may further improve the efficiency of the Chain-CAT. First, while we used the 2PL model, the 3PL model is also applicable. In going from 2PL to 3PL, the overall average RMSE values reported in Table 2 were found to improve by approximately 0.01. It is up to the analyst to decide which model to adopt, considering both the degree of improvement and the increased survey burden associated with the larger sample size required for the 3PL model. Second, although we used the length criterion as the stopping step, an alternative option is to use the precision criterion. This criterion administers as many items as necessary until pre-specified precision is obtained [23]. However, because the test length varies from respondent to respondent when the precision criterion is used, it would be inconvenient to administer the test in a classroom setting, so at-home administration may be preferable.
The algorithm for the use of collateral information could be improved. In the test step of the Chain-CAT algorithm, we directly used a given student’s proficiency estimate; namely, we chose the prior distribution as a normal distribution with for the second test. The other possibility is predicting the initial proficiency estimate of the th test by a regression model in a similar way to [34] or another session-series model using the given student’s proficiency estimates before the th test. Another option is to implement the algorithm utilizing the response time in addition to the response choice of the students [67]. This information might further improve the test efficiency.
Possible Practical Uses of Chain-CAT FCI
One practical application of Chain-CAT is its integration into ordinary short tests administered after university mechanics lectures. For example, each mini test could consist of eight questions: three problems related to the lecture content and five FCI questions presented in CAT format. Students would be graded only on the first three questions, while the latter five would not contribute to their grade. However, because all questions are presented consecutively on the LMS, it is expected that most students will complete the entire test to ensure their evaluation. This design is not only expected to evade declines in response rates but also allows instructors to achieve two complementary goals: The first three problems provide a measure of how well students have understood the lecture’s specific content. The five FCI questions assess students’ conceptual understanding of Newtonian mechanics in general at the time of the lecture. Throughout the semester, instructors can analyze the changes in conceptual understanding for each student.
As an example of how instructors might interpret Chain-CAT results, we could analyze the impact of specific instructional events, as suggested by Heckler and Sayre [17]. For instance, if providing feedback on homework is followed by a sharp increase in students’ conceptual understanding, this could serve as evidence of the effectiveness of the instructional method. Furthermore, one may find unexpected changes in the proficiency estimate, for example, decay, as observed in Ref. [17]. The state of understanding that appeared to have temporarily improved was actually unstable and can be interpreted as insufficient understanding. Researchers can use the Chain-CAT as an alternative tool to investigate the nature of student understanding.
Chain-CAT FCI would be particularly suitable for use in first-year university courses with high school physics as a prerequisite. In Japan, most first-year students who study mechanics at the university level have already encountered mechanics in high school, but their high school education tends to emphasize solving problems using formulas. University courses, especially those employing active learning methods such as tutorials, provide a more appropriate context for analyzing how students’ conceptual understanding develops.
@@
The following issues are needed to develop the feedback system based on Chain-CAT FCI. First, it is necessary to develop a method to analyze the progression of individual proficiency. Our simulation results showed a clear progression trend as shown in Fig. 2, because the statistics were calculated using 10 000 simulated responses at each session. However, for an individual respondent, the progression of proficiency fluctuates much more, which makes it difficult to understand the state of the student precisely. (As mentioned above, although the sample size is small, no significant fluctuations were observed in the surveys conducted with actual students. However, it remains possible that such fluctuations could appear in other contexts.) Therefore, we may need to estimate a student’s state based on the analysis of multiple test results, for example, by calculating an average across multiple sessions. It might be appropriate to group the proficiency estimates into several levels, as in Ref. [36].
There are many ways to provide feedback from Chain-CAT to respondents [68]. A simple score, the proficiency estimate would not be useful for students. It would be more useful, however, to group the proficiency estimates into levels, for example, simply dividing into four levels by setting thresholds. In Chain-CAT, students can receive feedback in the form of graphs showing their progress, as well as in the form of sentences with encouragement and/or guidance provided by generative artificial intelligence [69]. In considering what feedback to provide, student responses not only on the current test but also on prior tests should be considered. For example, a steadily increasing score across tests (indicating that all is well) might be met with automated feedback such as “I notice your scores are steadily improving. What strategies or habits do you think contributed most to this growth?” On the other hand, for a student with a stagnant score across several tests (indicating that additional effort is required) feedback might take the form of “I see your scores haven’t changed much yet. What parts of the test felt the hardest for you this time? Which ones felt more manageable?” In either case, this feedback will promote reflection, metacognition, and self-regulation for the students [70]. These tasks would be our future work.
Appendix A Item parameter estimates of the FCI using the 2PL model
| SE( | SE( | SE( | SE( | ||||||
| Item 1 | 1.113 | 0.064 | 0.961 | 0.059 | Item 16 | 1.231 | 0.066 | 0.710 | 0.048 |
| Item 2 | 0.844 | 0.052 | 0.425 | 0.056 | Item 17 | 1.453 | 0.071 | 0.525 | 0.041 |
| Item 3 | 0.923 | 0.068 | 1.915 | 0.123 | Item 18 | 1.568 | 0.074 | 0.069 | 0.036 |
| Item 4 | 1.064 | 0.058 | 0.586 | 0.051 | Item 19 | 1.059 | 0.059 | 0.533 | 0.049 |
| Item 5 | 1.453 | 0.070 | 0.225 | 0.038 | Item 20 | 1.233 | 0.063 | 0.343 | 0.042 |
| Item 6 | 0.900 | 0.060 | 1.365 | 0.088 | Item 21 | 0.980 | 0.055 | 0.471 | 0.052 |
| Item 7 | 0.942 | 0.058 | 0.906 | 0.064 | Item 22 | 1.145 | 0.060 | 0.125 | 0.043 |
| Item 8 | 1.201 | 0.069 | 1.102 | 0.060 | Item 23 | 1.351 | 0.066 | 0.276 | 0.040 |
| Item 9 | 1.140 | 0.061 | 0.444 | 0.046 | Item 24 | 1.254 | 0.067 | 0.769 | 0.049 |
| Item 10 | 1.571 | 0.096 | 1.458 | 0.065 | Item 25 | 1.893 | 0.086 | 0.156 | 0.033 |
| Item 11 | 1.327 | 0.065 | 0.229 | 0.040 | Item 26 | 1.966 | 0.093 | 0.567 | 0.035 |
| Item 12 | 1.156 | 0.079 | 1.777 | 0.098 | Item 27 | 0.881 | 0.054 | 0.592 | 0.058 |
| Item 13 | 2.916 | 0.139 | 0.061 | 0.028 | Item 28 | 1.188 | 0.064 | 0.634 | 0.047 |
| Item 14 | 1.135 | 0.059 | 0.096 | 0.043 | Item 29 | 0.182 | 0.049 | 6.645 | 1.799 |
| Item 15 | 0.591 | 0.047 | 0.716 | 0.087 | Item 30 | 2.011 | 0.093 | 0.435 | 0.034 |